# Disentangled Latent Energy-Based Style Translation: An Image-Level Structural MRI Harmonization Framework

### v1

@@ 4.1.2 Evaluation Tasks
Four tasks are performed in the experiments: (1) histogram and clustering comparison, (2) site classification, (3) brain tissue segmentation, and (4) site-specific MRI synthesis. The SRPBS is used in Tasks 1, 3, and 4 because traveling subjects are available. The OpenBHB is used in Task 2 and Task 4.
@@ 4.2 Experimental Results
4.2.1 Task 1: Histogram & Clustering Comparison
We first qualitatively evaluate our DLEST by comparing histogram and clustering of MRIs before harmonization (Baseline) and post-harmonization produced by each competing method in Fig. 2. The validation is performed on SRPBS since each subject has ground-truth scans on multiple acquisition sites. We treat the COI as the target domain due to its least intra-site variation, and harmonize the rest 10 sites/settings to the COI.
As observed in Fig. 2, before MRI harmonization (baseline), each site initially exhibits unique intensity distributions. Particularly, three prominent intensity peaks, corresponding to WM, GM, and CSF, are not aligned with the target domain. After harmonization, our DLEST framework demonstrates superior performance in aligning the histogram of the ten source sites with the target COI, particularly excelling in aligning the three main peaks. But the three SOTA deep learning methods (i.e., CycleGAN, StarGAN, and ImUnity) cannot well align source peaks with the target ones, despite showing good peak alignment across the 10 source sites. Our success may be attributed to the model’s ability to learn the latent data distribution of the entire target domain via EBM, while the three deep learning methods do not effectively account for intra-site variation within the target domain. On the other hand, the two non-learning methods cannot produce good performance. For example, the HM method, despite achieving reasonable peak alignment with the target, exhibits a non-smooth artifact in the post-harmonization histogram. This could be due to the significant intensity disparity between the source MRIs and the target/reference MRI (Li et al., 2021), caused by potential site effects.
For MRI clustering, we first employ a VGG network (Simonyan and Zisserman, 2014) to extract image features of the original MRIs and harmonized ones (via DLEST). Subsequently, we apply principle component analysis (with 20 principal components) to reduce the dimensionality of the extracted image features, and then utilize the t-SNE (Van der Maaten and Hinton, 2008) to reduce the dimensionality of these features and project them into a 2D space. Finally, we apply K-Means clustering algorithm to these features in this 2D space. The results in Fig. 3 show that image features are initially clustered into 11 distinct groups. After harmonization by DLEST, these clusters become less distinguishable, implying that our method effectively eliminates site-related variations in harmonized MRIs.
Figure 3: Clustering of 11 sites across 9 subjects from SRPBS, (left) before and (right) after harmonization by DLEST.
@@ 4.2.2 Task 2: Site Classification
In this task, we conduct a quantitative evaluation to determine the effectiveness of a harmonization method in eliminating site-related, non-biological variations. We begin by selecting a site (ID: 17) from the OpenBHB dataset with minimal intra-site variation as our target domain. We then harmonize MRIs from the rest 57 source sites to this target domain using each method. Following this, we train a site classifier on the OpenBHB training set (with $3,227$ MRIs) and validate it on the validation set (with $757$ MRIs). The site classifier is implemented based on the ResNet18 (He et al., 2016) architecture. Its performance on raw MRI data is recorded as the Baseline. Five metrics are used: balanced accuracy (BACC), accuracy (ACC), area under the ROC curve (AUC), F1 score (F1), and sensitivity (SEN). The results of site classification are reported in Table 1, where a lower value signifies a method’s better capability to eliminate site-related variations, thereby making it challenging for the site classifier to learn site-related features that could differentiate each site.
@@
Figure 4: Segmentation results of CSF, GM, and WM on SRPBS in terms of (left) Dice coefficient and (right) Jaccard index. Each U-Net is trained on the target site COI and validated the source site HUH from SRPBS.
Figure 5: Sample U-Net segmentation of brain tissues (CSF, GM, WM) using unharmonized data (Baseline) and data harmonized by CycleGAN (top competing method) and DLEST (ours). White pixels denote the correct segmentation, red pixels denote under-segmentation, and blue pixels denote over-segmentation.
Figure 6: Synthesized MRI with site-specific styles of 3 sites from SRPBS (top 3 rows) and 1 site from OpenBHB (bottom row). The synthetic MRIs have diverse anatomical structures with consistent image styles that match the reference MRI from a specific site.
@@ 4.2.3 Task 3: Brain Tissue Segmentation
This experiment is designed to assess the effectiveness of the DLEST on a downstream brain tissue (i.e., WM, GM, and CSF) segmentation task. We begin by selecting a source site (i.e., HUH) and a target site (i.e., COI) within the SRPBS dataset and harmonize the source MRIs to the target domain using DLEST and each competing method. Following this, a U-Net (Ronneberger et al., 2015) segmentation model is trained on MRIs from the target site (COI) and directly validated on the source site (HUH). As our primary objective is to assess the impact of each harmonization method on tissue segmentation, the quality of the ground-truth segmentation is not our main concern. We thus utilize the segmentation masks obtained using the FAST pipeline in FSL as our ground truth to provide a fair basis for comparing each method. For each harmonization method, the U-Net is trained for 100 epochs, with the mean and standard deviation of the validation Dice coefficient and the Jaccard index reported for the final 40 epochs.
Quantitatively, as shown in Fig. 4, our DLEST method consistently surpasses all non-learning methods (i.e., HM and SSIMH), suggesting that the latent space style translation is more effective than merely aligning the global intensity distribution. Compared with three SOTA methods, our DLEST consistently delivers superior performance in terms of average segmentation results and specifically in the segmentation of CSF and GM. While CycleGAN slightly outperforms DLEST in WM segmentation, our DLEST exhibits a significantly lower standard deviation. This indicates that the DLEST produces a more stable performance compared to CycleGAN.
@@ 4.2.4 Task 4: MRI Synthesis
This task aims to demonstrate our framework’s ability to synthesize brain MRIs with site-specific styles. We select three sites from the SRPBS dataset and one site from the OpenBHB dataset for illustration. We evaluate the quality of the synthesis qualitatively via sample visualization and quantitatively through the structural similarity index measurement (SSIM) (Wang et al., 2004), peak signal-to-noise ratio (PSNR), and root mean square error (RMSE). As shown in Fig. 6, our DLEST is able to synthesize new brain MRIs with diverse brain shapes and anatomical structures while preserving the site-specific styles that match the given reference MRI scans. This latent space MRI synthesis is cost-efficient and helpful in training non-disease specific learning models that require large quantities of MRI samples while the actual MRI data is limited, such as brain segmentation models.
Table 2: Quantitative result of the image-level metrics. Original denotes mean and standard deviation between every pair of original images; and Synthesized denotes mean and standard deviation between every pair of original and synthesized images.
@@
COI Original $0.829$$\pm$$0.025$ $21.590$$\pm$$1.718$ $0.085$$\pm$$0.013$
Synthesized $0.828$$\pm$$0.017$ $20.740$$\pm$$1.485$ $0.093$$\pm$$0.016$
YC1 Original $0.838$$\pm$$0.028$ $19.883$$\pm$$2.236$ $0.104$$\pm$$0.022$
Synthesized $0.830$$\pm$$0.019$ $19.723$$\pm$$1.456$ $0.105$$\pm$$0.018$
UTO Original $0.840$$\pm$$0.031$ $21.387$$\pm$$2.323$ $0.088$$\pm$$0.018$
Synthesized $0.833$$\pm$$0.017$ $21.405$$\pm$$1.372$ $0.086$$\pm$$0.014$
Site 17 Original $0.800$$\pm$$0.025$ $18.936$$\pm$$1.541$ $0.115$$\pm$$0.016$
Synthesized $0.788$$\pm$$0.015$ $18.421$$\pm$$0.951$ $0.121$$\pm$$0.013$
For quantitative evaluation, we report the mean and the standard deviation of image-level metrics for synthesized MRIs at each site in Table 2. Metrics under the Original category are derived from comparisons between each original image and every other original image within the same site. The Synthesized category metrics are obtained by comparing each synthesized image with all original images from the same site. The marginal differences observed across all metrics between Original and Synthesized image pairs suggest that our DLEST framework can effectively generate synthetic MRIs that maintain consistent styles and quality comparable to the original images.
@@ 5.1 Ablation Study
We investigate three downgraded variants of the proposed DLEST framework, namely (1) DLEST-PIX, which excludes pixel-wise loss in Eq. (5) during the SIG module training, (2) DLEST-CON, which omits the latent content loss in Eq. (10), and (3) DLEST-CYC, which does not incorporate the latent cycle-consistency loss in Eq. (10) when training the SST module for style translation. We train DLEST and its three variants on the SRPBS traveling subject dataset and evaluate their harmonization results using two image quality metrics: SSIM and PSNR. As indicated in Fig. 7 (a), while all ablations negatively impact the result, DLEST-PIX particularly yields the worst results, compared with its two counterparts. This implies that the pixel-wise, latent content and cycle-consistency losses are necessary to ensure a satisfying harmonization outcome, and especially, the pixel-wise loss plays a crucial role in preserving image content in DLEST.
@@ 5.2 Influence of Hyperparameters
We conduct several rounds of hyperparameter tuning to examine the influence of hyperparameters $\alpha$ and $\beta$ in Eq. (10) on the harmonization effect. This involved varying one parameter while maintaining the other constant. The results in Fig. 7 (b)-(c) suggest that the model’s performance improves when the latent content loss has more weight (i.e., with a larger $\alpha$) in the overall loss function. Additionally, the performance of DLEST appears to be less sensitive to variations in $\beta$.
@@ 5.3 Influence of Content Loss
Following (Zhao and Chen, 2021), our implementation uses a latent content loss in Eq. (8) to minimize the change of anatomical structure during style translation. This implicit regularization in latent space may cause slight alterations of anatomical details. To this end, we compare our method with its variant DLEST+CS that utilizes a separate style and content loss (Gatys et al., 2016) to explicitly enforce content preservation while allowing the style to be translated freely. In DLEST+CS, a pre-trained VGG19 (Simonyan and Zisserman, 2014) is used to extract feature maps of given MRI slices. The content loss is calculated as the $l_{2}$ distance between the extracted feature maps of original MRI and the translated MRI in specific layers. The style loss is computed as the weighted average of $l_{2}$ distance between Gram matrices (Gatys et al., 2016) of the target and translated MRIs in each layer. The Gram matrix is a correlation matrix between each feature map, where each entry $\mathcal{G}_{ij}$ is the inner product of vectorized feature maps $i$ and $j$. By including feature correlations of multiple layers, we obtain a multi-scale representation of the input image, which captures its texture characteristics without content information (Gatys et al., 2016). As shown in Fig. 7 (a), compared with DLEST, its variant DLEST+CS achieves slight improvement in terms of SSIM, implying enhanced content preservation. However, the improvement of DLEST+CS is accompanied by a significant increase in computation cost, as the calculation of style and content loss in ambient space, which requires decoding latent codes into image slices within EBM.
Figure 7: Results of (a) DLEST variants, DLEST with different values of (b) $\alpha$ and (c) $\beta$, and (d) different ratios of style/content loss.
5.4 Comparison with Feature-Level Harmonization
Although the primary motivation behind proposing an image-level harmonization approach is its adaptability to a variety of downstream tasks, it remains crucial to compare our DLEST framework with the prevalent feature-level harmonization method, ComBat (Johnson et al., 2007). This comparison generally follows the evaluation scheme outlined in the OpenBHB Challenge (Dufumier et al., 2022) to ensure a fair assessment of both methodologies. Specifically, we employ the official Python implementation of ComBat (Fortin, 2023; Johnson et al., 2007) to harmonize a collection of seven pre-extracted MRI features. These features include cortical thickness (along with standard deviation), GM volume, surface area, integrated mean (and Gaussian) curvature, and intrinsic curvature index for each subject in the training set. Subsequently, we train a linear logistic regression model (Pedregosa et al., 2011) on these harmonized features and evaluate its multi-class site classification performance on the validation set. Similarly, for the image-level approach, we utilize a VGG19 feature extractor (with weights frozen) to directly extract image features from the harmonized MRIs, and train a comparable linear logistic regression model for site classification. We also compare two baselines, named Baseline-F and Baseline-I that correspond to the original MRI features and images, respectively. The accuracy and balanced accuracy, and F1 score are reported, with lower performance indicating more effective harmonization. As shown in Table 3, both the original MRI scans and extracted MRI features are characterized by a significant amount of site-related variations. ComBat performs better than Baseline-F, indicating its efficacy in feature harmonization. However, it only addresses the site effect on the pre-extracted features, rather than the entire image level, limiting its utility in broader applications. In contrast, our method successfully eliminates the majority of the site-related variations, resulting in the lowest BACC for the linear model.
Table 3: Site classification result of linear model where ‘*’ denotes the difference between a specific method and DLEST is statistically significant (p $<$ 0.05 via paired t-test).
[…]
Baseline-F $0.703\pm 0.019^{*}$ $0.762\pm 0.008^{*}$ $0.739\pm 0.020^{*}$
Baseline-I $0.311\pm 0.087^{*}$ $0.505\pm 0.086^{*}$ $0.451\pm 0.083^{*}$
ComBat $0.586\pm 0.158^{*}$ $0.566\pm 0.213^{*}$ $0.525\pm 0.230^{*}$
CycleGAN $0.134\pm 0.029$ $0.392\pm 0.060$ $0.303\pm 0.046^{*}$
StarGAN $0.178\pm 0.053^{*}$ $0.430\pm 0.112$ $0.352\pm 0.088^{*}$
ImUnity $0.196\pm 0.043^{*}$ $0.406\pm 0.107$ $0.350\pm 0.080^{*}$
DLEST (Ours) $0.130\pm 0.010$ $0.407\pm 0.013$ $0.321\pm 0.010$

### v2

@@ 4.1.2 Evaluation Tasks
Four tasks are performed, including (1) histogram comparison and feature visualization, (2) site classification, (3) brain tissue segmentation, and (4) site-specific MRI synthesis. The SRPBS is used in Tasks 1, 3, and 4 because traveling subjects are available. The OpenBHB is used in Task 2 and Task 4.
@@ 4.2 Experimental Results
4.2.1 Task 1: Histogram Comparison & Feature Visualization
We first qualitatively evaluate our DLEST by comparing histograms and extracted features of MRIs before harmonization (Baseline) and post-harmonization produced by each competing method in Fig. 2. The validation is performed on SRPBS since each subject has ground-truth scans on multiple acquisition sites. To select a target site, we compare intra-site variations across each site in the SRPBS dataset by computing the mean peak signal-to-noise ratio (PSNR) for all image pairs within a given site. Given that the SRPBS encompasses all traveling subjects, each site shares an identical subject cohort, thereby containing consistent anatomical information. Consequently, a site yielding a higher intra-site PSNR score suggests a greater similarity in image style within that particular site. We treat COI as the target domain due to its low intra-site variation and harmonize the remaining 10 sites/settings to the COI. The influence of target site selection is elaborated in Section 5.4.
As observed in Fig. 2, before MRI harmonization (baseline), each site initially exhibits unique intensity distributions. In particular, three prominent intensity peaks, WM, GM, and CSF, are not aligned with the target domain. After harmonization, our DLEST framework demonstrates superior performance in aligning the histogram of the ten source sites with the target COI, particularly excelling in aligning the three main peaks. However, the three SOTA deep learning methods (i.e., CycleGAN, StarGAN, and ImUnity) cannot well align source peaks with the target ones, despite showing good peak alignment across the 10 source sites. Our success may be attributed to the model’s ability to learn the latent data distribution of the entire target domain via EBM, while the three deep learning methods do not effectively account for intra-site variation within the target domain. On the other hand, the two non-learning methods cannot produce good performance. For example, the HM method, despite achieving reasonable peak alignment with the target, exhibits a non-smooth artifact in the post-harmonization histogram. This could be due to the significant intensity disparity between the source MRIs and the target/reference MRI [22], caused by potential site effects.
For feature visualization, we use a pre-trained ResNet18 [15] model to extract features of the original MRIs and harmonized ones (via DLEST) across all sites and subjects. We then apply principal component analysis (with 100 principal components) to reduce the dimensionality of the extracted image features, and then utilize the t-SNE [24] to reduce the dimensionality of these features further and project them into a 2D space. Finally, we color-code each site with a unique color. The visualization results in Fig. 3 show that the extracted image features are initially clustered into 11 distinct groups. After harmonization by DLEST, these clusters become less distinguishable, implying that our method effectively eliminates site-related variations.
Figure 3: Visualization of MRI features of 11 sites across 9 subjects from SRPBS, with MRI (a) before and (b) after harmonization by DLEST. Each color denotes a specific site, while each point denotes extracted features from an MRI slice of a specific subject.
@@ 4.2.2 Task 2: Site Classification
In this task, we conduct a quantitative evaluation to determine the effectiveness of a harmonization method in eliminating site-related, non-biological variations. We begin by selecting a site (ID: 17) from OpenBHB with minimal intra-site variation as our target site. We then harmonize MRIs from the rest 57 source sites to this target domain using each method. Following this, we train a site classifier on OpenBHB training set (with $3,227$ MRIs) and validate it on the validation set (with $757$ MRIs). The site classifier is implemented based on the ResNet18 [15] architecture. Its performance on raw MRI data is recorded as the Baseline. Five metrics are used: balanced accuracy (BACC), accuracy (ACC), area under the ROC curve (AUC), F1 score (F1), and sensitivity (SEN). The results of site classification are reported in Table 1, where a lower value signifies a method’s better capability to eliminate site-related variations, thereby making it challenging for the site classifier to learn site-related features that could differentiate each site.
@@
Figure 4: Segmentation results of CSF, GM, and WM on SRPBS in terms of (left) Dice coefficient and (right) Jaccard index. Each U-Net is trained on the target site COI and validated on the source site HUH from SRPBS.
Figure 5: Segmentation of brain tissues (i.e., CSF, GM, WM) using unharmonized MRI data (Baseline) and data harmonized by CycleGAN (top competing method) and DLEST (ours). White pixels denote the correct segmentation, red pixels denote under-segmentation, and blue pixels denote over-segmentation.
@@
Figure 6: Synthesized MRI with site-specific styles of 3 sites from SRPBS (top 3 rows) and 1 site from OpenBHB (bottom row). The synthesized MRIs have diverse anatomical structures with consistent image styles that match the reference MRI from a specific site.
@@ 4.2.3 Task 3: Brain Tissue Segmentation
This experiment is designed to assess the effectiveness of the DLEST on a downstream brain tissue (i.e., WM, GM, and CSF) segmentation task. We begin by selecting a source site (i.e., HUH) and a target site (i.e., COI) within the SRPBS dataset and harmonize the source MRIs to the target domain using DLEST and each competing method. Similar to [23], a U-Net [32] segmentation model is trained on MRIs from the target site (COI) and directly validated on the source site (HUH). As our primary objective is to assess the impact of each harmonization method on tissue segmentation, the quality of the ground-truth segmentation is not our main concern. We thus utilize the segmentation masks obtained using the FAST pipeline in FSL as our ground truth for UNet training, providing a fair basis for comparing each method. For each harmonization method, the U-Net is trained for 100 epochs, with the mean and standard deviation of the validation Dice coefficient and the Jaccard index reported for the final 40 epochs.
Figure 4 quantitatively illustrates that despite the source and target sites sharing the same subject cohort, the direct application of a UNet model trained on the site COI to raw data from site HUH results in less than optimal segmentation. This suggests that site-related variations hinder the generalizability of the segmentation model. Our DLEST method consistently surpasses all non-learning methods (i.e., HM and SSIMH), suggesting that the latent space style translation is more effective than merely aligning the global intensity distribution. Compared with the three SOTA methods, our DLEST consistently delivers improved segmentation performance in terms of average segmentation results specifically in the segmentation of CSF and GM. While CycleGAN slightly outperforms DLEST in WM segmentation, our DLEST exhibits a significantly lower standard deviation. This indicates that the DLEST produces a more stable performance compared with CycleGAN.
@@ 4.2.4 Task 4: MRI Synthesis
This task aims to demonstrate our framework’s ability to synthesize brain MRIs with site-specific styles. We select three sites from SRPBS and one site from OpenBHB for illustration. We evaluate the synthesized MRIs qualitatively via sample visualization and quantitatively using structural similarity index measurement (SSIM) [43], peak signal-to-noise ratio (PSNR), root mean square error (RMSE), and intensity Pearson correlation coefficient (PCC). In our analysis, we apply a binary brain mask to eliminate any influence of the background pixels when calculating each metric. As shown in Fig. 6, our DLEST is able to synthesize new brain MRIs with diverse brain shapes and anatomical structures while preserving the site-specific styles that match the given reference MRI scans. This latent space MRI synthesis is cost-efficient and helpful in training non-disease-specific learning models that require large quantities of MRI samples while the actual MRI data is limited, such as brain segmentation models.
Table 2: Quantitative result of the image-level metrics. The term “Original MRI” denotes mean and standard deviation between every pair of original images; and “Synthesized MRI” denotes mean and standard deviation between every pair of original and synthesized images.
@@
COI Original MRI $0.900$$\pm$$0.025$ $27.206$$\pm$$2.266$ $0.045$$\pm$$0.010$ $0.916$$\pm$$0.037$
Synthesized MRI $0.901$$\pm$$0.019$ $26.563$$\pm$$1.817$ $0.048$$\pm$$0.010$ $0.925$$\pm$$0.028$
YC1 Original MRI $0.900$$\pm$$0.029$ $25.386$$\pm$$2.672$ $0.056$$\pm$$0.015$ $0.933$$\pm$$0.033$
Synthesized MRI $0.903$$\pm$$0.019$ $25.584$$\pm$$1.852$ $0.054$$\pm$$0.011$ $0.942$$\pm$$0.023$
UTO Original MRI $0.904$$\pm$$0.026$ $27.041$$\pm$$2.572$ $0.046$$\pm$$0.011$ $0.905$$\pm$$0.041$
Synthesized MRI $0.905$$\pm$$0.018$ $27.249$$\pm$$1.680$ $0.044$$\pm$$0.009$ $0.914$$\pm$$0.029$
Site 17 Original MRI $0.879$$\pm$$0.022$ $25.176$$\pm$$1.892$ $0.056$$\pm$$0.011$ $0.899$$\pm$$0.036$
Synthesized MRI $0.873$$\pm$$0.015$ $24.918$$\pm$$1.270$ $0.057$$\pm$$0.009$ $0.898$$\pm$$0.031$
For quantitative evaluation, we report the mean and the standard deviation of image-level metrics for synthesized MRIs at each site in Table 2. Results under the Original category are derived from comparisons between each original image and every other original image within the same site. The Synthesized category results are obtained by comparing each synthesized image with all original images from the same site. The negligible disparity across all four metrics between the original-original and original-synthesized image pairs implies that the synthesized images uphold the same quality as their original counterparts. Moreover, the slightly increased PCC observed in most sites indicates that the synthesized MRI slices exhibit a consistent style and fewer variations when compared to the original MRIs from the respective sites.
Figure 7: MRI quality assessment results achieved by DLEST and its variants on traveling subjects from SRPBS.
@@ 5.1 Ablation Study
We investigate three downgraded variants of the proposed DLEST framework, namely (1) DLEST-PIX, which excludes pixel-wise loss in Eq. (5) during the SIG module training, (2) DLEST-CON, which omits the latent content loss in Eq. (10), and (3) DLEST-CYC, which does not incorporate the latent cycle-consistency loss in Eq. (10) when training the SST module for style translation. We train DLEST and its three variants on the SRPBS traveling subject dataset and evaluate their harmonization results using three image quality metrics: SSIM, PSNR, and PCC. As indicated in Fig. 7, while all ablations negatively impact the result, DLEST-PIX particularly yields the worst results, compared with its two counterparts. This implies that the pixel-wise, latent content and cycle-consistency losses are necessary to ensure a satisfying harmonization outcome, and especially, the pixel-wise loss plays a crucial role in preserving image content in DLEST.
Figure 8: Results of DLEST with different values of (a) $\alpha$ and (b) $\beta$ on the SRPBS traveling subject dataset.
@@ 5.2 Influence of Hyperparameters
We conduct several rounds of hyperparameter tuning to examine the influence of hyperparameters $\alpha$ and $\beta$ in Eq. (10) on the harmonization effect. This involves varying one parameter while maintaining the other constant. The results in Fig. 8 suggest that the model’s performance improves when the latent content loss has a balanced weight (i.e., $\alpha=1$). Additionally, the performance of DLEST appears to be less sensitive to variations in $\beta$ with optimal performance at $\beta=100$.
Figure 9: (a) Comparison of DLEST and DLEST+CS, and (b) results of DLEST+CS with different style/content ratios on the SRPBS traveling subject dataset.
@@ 5.3 Influence of Content Loss
The proposed DLEST uses a latent content loss in Eq. (8) to constrain the change of latent information while minimizing the $\mathcal{L}_{EBM}$ during style translation. This implicit regularization in latent space may cause slight alterations in anatomical details. To this end, we compare our method with its variant DLEST+CS that utilizes a separate style and content loss [11] to explicitly enforce content preservation while allowing the style to be translated freely. In DLEST+CS, a pre-trained VGG19 [35] is used to extract feature maps of given MRI slices. The content loss is calculated as the $l_{2}$ distance between the extracted feature maps of original MRI and the translated MRI in specific layers. The style loss is computed as the weighted average of $l_{2}$ distance between Gram matrices [11] of the target and translated MRIs in each layer. The Gram matrix is a correlation matrix between each feature map, where each entry $\mathcal{G}_{ij}$ is the inner product of vectorized feature maps $i$ and $j$. By including feature correlations of multiple layers, we obtain a multi-scale representation of the input image, which captures its texture characteristics without content information [11].
As shown in Fig. 9 (a), compared with DLEST, its variant DLEST+CS achieves slight improvement across all metrics, implying enhanced content preservation and style translation. We also investigate the influence of different ratios of style and content loss as suggested in [11]. The results in Fig. 9 (b) indicate that an emphasis on content loss (lower style/content loss ratio) leads to better harmonization performance. Despite the enhancements by DLEST+CS, it’s important to note that these come with a substantial rise in computational expense. This is primarily due to the computation of style and content loss in the ambient image space. Consequently, after each iteration of Stochastic Gradient Langevin Dynamics (SGLD), the latent code must be decoded. Furthermore, this process necessitates the use of an additional pre-trained feature extractor, adding to the complexity and computational demand.
Figure 10: Histogram of harmonized MRIs using four different target sites (i.e., YC1, UTO, HUH, and HKH) on SRPBS. Given a specific target site, MRIs from the remaining 10 sites are used as source data. The PSNR value denotes the intra-site variation of each target site.
5.4 Influence of Target Site
We further examine the impact of target site selection on harmonization using four different target sites (i.e., YC1, UTO, HUH, and HKH) in the SRPBS dataset. Similar to the main experiments, given a selected target site, we harmonize the remaining 10 sites as the source domain. Notably, only the SST module requires fine-tuning with each target change, typically requiring about 100 2D MRI slices from the target domain. This process incurs minimal computational cost, as detailed in Section 5.5. The histogram results of the harmonized MRIs, depicted in Fig. 10, underscore the influence of target site selection on the harmonization outcome. When a target site with lower intra-site variations (higher PSNR), such as site HKH, is used, the harmonized histograms across all sites align more effectively. Conversely, a target site with a lower PSNR, such as site HUH, may result in less optimal histogram alignment. This observation suggests that a target site with excessive variations could hinder the model’s ability to accurately capture the latent data distribution that characterizes the target style.
Table 3: Results of time and computation cost. For DLEST, “$a+b$” denotes the number for SIG $+$ SST. The term “Test time” is the cost of generalizing on an unseen dataset. M: Million; GMac: Giga multiply-accumulate operations; H: Hour.
Method ##\##Parameters (M) FLOPs (GMac) Training Time (H) Test Time (H) Need Re-training
CycleGAN $28.3$ $1794.1$ $29.1$ $8.0$ $Yes$
StyleGAN $161.3$ $4865.3$ $53.8$ $9.1$ $Yes$
ImUnity $671.0$ $246.2$ $25.3$ $1.5$ $No$
DLEST (Ours) $21.7$+$5.3$ $2.1$+$0.3$ $20.0$+$0.4$ $0.4$ $No$